OpenStudy (anonymous):

Evaluate. 7 × [8 + (16 – 9)] ÷ 5

OpenStudy (oaktree):

So we're going to start off with the innermost parenthesis, right? So what are we going to do first?

OpenStudy (anonymous):

i don't know ANYTHING about Evaluating

OpenStudy (anonymous):

im not that good at it

OpenStudy (anonymous):

it will be 21

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can u maybe explain?

OpenStudy (anonymous):

The Order of Operations is: Parentheses Exponents Multiplication Division Addition Subtraction Your problem is 7x[8+(16-9)]/5 According to the Order of Operstions, we start with the innermost parentheses, which would be 16-9. 16-9=7 Our new equation looks like this: 7x[8+7]/5 We still have a pair of parentheses, so we solve that next. [8+7}=15. Our new equation looks like this: 7x15/5. According to the Order of Operations, we now multiply 7x15=105 Our new eqution looks like this:105/5 105/5=21, so that is our answer.